OpenStudy (anonymous):

If the odds of winning a raffle are 12:258, what is the probability of winning? 12/270 12/258 258/270 258/12

OpenStudy (anonymous):

would that be 258/12

OpenStudy (anonymous):

answer is in the question

OpenStudy (anonymous):

replace : with /

OpenStudy (anonymous):

12/258

OpenStudy (anonymous):

is this the right way....had it backwards

OpenStudy (anonymous):

ya

hartnn (hartnn):

if p is the probability,the odds of winning is p/(1-p)

hartnn (hartnn):

258/12 is the incorrect answer you will get correct answer by solving this \[\frac{p}{1-p}=\frac{12}{258}\] do u know how to solve it?

OpenStudy (anonymous):

so that would be 12(258+12)=12/270 is that close

hartnn (hartnn):

yup that is correct :)